### v1

# Establishing Minimum Elements for Effective Vulnerability Management in AI Software

Abstract
In the rapidly evolving field of artificial intelligence (AI), the identification, documentation, and mitigation of vulnerabilities are paramount to ensuring robust and secure systems. This paper discusses the minimum elements for AI vulnerability management and the establishment of an Artificial Intelligence Vulnerability Database (AIVD). It presents standardized formats and protocols for disclosing, analyzing, cataloging, and documenting AI vulnerabilities. It discusses how such an AI incident database must extend beyond the traditional scope of vulnerabilities by focusing on the unique aspects of AI systems. Additionally, this paper highlights challenges and gaps in AI Vulnerability Management, including the need for new severity scores, weakness enumeration systems, and comprehensive mitigation strategies specifically designed to address the multifaceted nature of AI vulnerabilities.
Index Terms:
AI Incidents, AIVD, AI Vulnerability Management, AI Bill of Materials, AIBOM
I Introduction
As artificial intelligence (AI) technologies permeate various sectors including healthcare [1], finance [2], transportation [3], industry[4, 5], and security [6, 7, 8], their transformative impact is undeniable. However, this rapid integration introduces complex vulnerabilities that may compromise the systems it is intended to enhance. With AI becoming a foundational component across various sectors, there is an urgent need to systematically identify, evaluate, and mitigate AI-specific vulnerabilities. Therefore, Vulnerability Management in AI systems is a critical process that encompasses the identification, assessment, public disclosure, and remediation of security vulnerabilities within the systems and components. This practice is essential due to the unique and complex nature of AI software, which often includes data-driven decision-making processes that are opaque and challenging to secure.
The establishment of a robust vulnerability management framework for AI systems underscores the need for centralized information sharing, where AI vulnerabilities can be reported, cataloged, and communicated with all stakeholders. This centralization ensures that information regarding new and emerging threats is quickly and widely disseminated, allowing for more rapid responses across the industry. This demands a global AI vulnerability database to serve as a central repository for information on identified vulnerabilities specific to AI systems, providing crucial support for organizations in protecting their AI technologies against security threats. While the National Vulnerability Database (NVD) [9], managed by the National Institute of Standards and Technology (NIST), offers a proven framework for cataloging software vulnerabilities, the entire vulnerability management ecosystem, is not designed to scale to support AI vulnerabilities and incidents.
AI vulnerabilities differ significantly from traditional software vulnerabilities due to the adaptive and often data-driven nature of AI systems. Unlike conventional software, where vulnerabilities typically arise from coding errors or system misconfigurations, AI vulnerabilities often stem from the training data, model architecture, or the learning algorithms themselves. For instance, adversarial attacks[10], where malicious inputs are designed to fool AI models into making incorrect decisions, highlight vulnerabilities unique to AI systems. These vulnerabilities can be exploited to cause AI systems to behave unpredictably or to make decisions that benefit attackers. Understanding these vulnerabilities requires a deep dive into the underlying algorithms and data, making the process of identifying and mitigating these vulnerabilities complex and challenging.
Given the unique nature of AI vulnerabilities, there is a clear need for a specialized repository that can serve as a central hub for information on AI-specific vulnerabilities. The proposed Artificial Intelligence Vulnerability Database (AIVD) would function similarly to the NVD but would be specifically tailored to catalog and disseminate information about vulnerabilities unique to AI systems. In this paper, we discuss the minimum elements for AI vulnerability management and establishing an Artificial Intelligence Vulnerability Database. We introduce standardized formats and protocols for disclosing, analyzing, cataloging, and measuring AI vulnerabilities. We discuss how such AI incident database must extend beyond the traditional scope of the NVD by focusing on the unique aspects of AI systems. It needs to catalog issues such as biases in training data, flaws in algorithmic design, and vulnerabilities specific to AI deployment environments. Additionally, we highlight challenges and gaps in AI vulnerability management such as the need for new severity scores, weakness enumeration systems, and comprehensive mitigation strategies, specifically designed to address the multifaceted nature of AI vulnerabilities. These systems, dependent on data quality and algorithmic complexity, require adaptive approaches due to their dynamic, non-deterministic nature. SectionVI further discusses the gaps and challenges in AI vulnerability management.
Item Field Description
1 AI-CVE ID This is a unique identifier assigned to each recorded vulnerability within the AIVD to ensure easy tracking and referencing.
2 AI Model Details Specific details about the AI model or system affected (e.g., model type, version).
3 Weakness Type Identifies the type of vulnerability found in the AI system, providing details on the general category or the specific mechanism by which the system is compromised. Similar to CWE in NVD database, each vulnerability will be assigned with an AI-CWE.
4 Root Cause Explains the underlying factors contributing to the vulnerability. This root cause could be due to issues such as inherent vulnerabilities within the model, inadequate resiliency and robustness safeguards, and flaws in data.
5 Impact How the vulnerability is affecting the product.
6 Severity Scores Adapted to reflect the unique impacts and severity considerations of AI vulnerabilities. Determining the severity score is crucial because it quantifies the severity of vulnerabilities, guiding prioritization and response strategies.
7 Affected Software & Products Applications susceptible to the vulnerability, as it allows organizations to precisely target their remediation efforts, manage risks effectively, and ensure compliance with security standards.
8 Exploitability It explains how a vulnerability can be exploited, focusing on the technical complexity of the attack, the level of privileges necessary, and the specific actions required by users to exploit the vulnerability.
9 Description A thorough explanation of the vulnerability, or directst to an external source for a more comprehensive description of the vulnerability.
10 Mitigation Outlines specific steps and strategies to effectively address and neutralize a vulnerability within an AI system, taking into account the unique characteristics and behaviors of AI technologies. It includes detailed recommendations for patching flaws, strengthening security protocols, and updating AI models to resist attacks.
11 References Relevant documents, papers, and reports of the vulnerability.
12 Report Date The date on which the vulnerability was first identified or reported.
13 Reported By The individual or organization that first reported the vulnerability.
14 Vendor Identifies the developer or the supplier of the AI system or component which is affected by the vulnerability.
15 Status Current status of the vulnerability (e.g., unresolved, patched).
TABLE I: Details of AI Vulnerability Assessment
II Minimum Elements for Defining AI Vulnerabilities
This section describes the Minimum Elements (MEs) essential for capturing and describing AI vulnerabilities effectively. Each ME represents a critical piece of information that contributes to the comprehensive profiling and understanding of vulnerabilities within AI systems. These elements facilitate structured data collection, enabling systematic tracking, analysis, and mitigation of vulnerabilities, and equip stakeholders with necessary details to effectively address vulnerabilities from identification through resolution. Each ME is strategically defined to encapsulate crucial information regarding the vulnerabilities it describes. This approach not only standardizes vulnerability reporting across various platforms and industries but also ensures that all stakeholders—from developers and security professionals to regulatory bodies—have access to the same comprehensive data set. The structured data collection enabled by these elements allows organizations to systematically track vulnerabilities from initial identification to final resolution. This process includes detailed analysis stages where vulnerabilities are assessed for impact and exploitability, allowing for informed prioritization in mitigation efforts.
Furthermore, these elements serve as the backbone for creating databases and registries that stakeholders can reference to ensure consistency and accuracy in vulnerability management. By defining these minimum required data points, the AI vulnerability disclosure process becomes more efficient, transparent, and scalable, accommodating the dynamic nature of AI technologies and the complex ecosystems in which they operate. This standardized approach not only streamlines the mitigation process but also enhances the security posture of AI systems by providing clear and actionable information that can be rapidly deployed to counter potential threats. These minimum elements are described in the Table I. Some of these minimum elements require further advancements in foundational knowledge and the creation of new frameworks, and catalogs to describe AI vulnerabilities.
II-A Weakness Type
Despite their sophisticated capabilities, AI systems are uniquely vulnerable to certain types of weaknesses that are seldom encountered in traditional software systems. These vulnerabilities stem from inherent characteristics of AI, such as heavy reliance on expansive datasets and complex decision-making processes that lack transparency. While resources like CWE exist to catalog vulnerabilities in traditional software systems, there is currently no equivalent resource specifically tailored for AI. Moreover, the CWE framework cannot fully capture the unique weaknesses of AI systems, such as those introduced by their reliance on data and algorithmic complexity. AI’s dependency on data quality and volume means that any corruption or bias in the data can significantly impact performance and decision accuracy. Furthermore, the black-box nature of many AI algorithms complicates the detection and understanding of vulnerabilities, making it challenging to develop robust AI systems.
Building on the need to address the unique vulnerabilities of AI systems, we propose the development of an AI-specific Common Weakness Enumeration (AI-CWE). AI-CWE will comprehensively catalog and enumerate AI-specific weaknesses, thereby facilitating a more structured approach to identifying these vulnerabilities. To structure the AI-CWE, we draw upon the AI/ML ATT&CK framework[11], ensuring that each identified weakness is linked to a specific deficiency in the system’s mitigation strategies. For instance, the four main classes of weaknesses are as follows:
  1. 1.
    Insufficient validation mechanisms that allow malicious samples to bypass security checks and infiltrate the system.
  2. 2.
    Data handling processes lack robust filtering and normalization capabilities, leading to susceptibility to noise and perturbations that compromise data integrity.
  3. 3.
    The learning algorithms are vulnerable to perturbations due to inadequate resilience and adaptability in handling crafted inputs designed to mislead or corrupt the learning process.
  4. 4.
    Privacy safeguards are deficient or absent, exposing the data to unauthorized access and manipulation due to inadequate encryption, anonymization, or access control configurations.
Each of these weaknesses will be further broken down into more detailed categories for analysis and mitigation. This categorization allows users to better understand the specific type of weakness within their AI system, enabling them to reason about the root cause of the weakness and employ more targeted and effective mitigation strategies. Tables II and III present two samples of AI-CWEs and the components to consider in describing them.
AI-CWE-ID AI-CWE-100
AI-CWE-Name Inadequate Input Filtering
Description The data handling processes lack robust filtering and normalization capabilities, leading to susceptibility to noise and perturbations which compromise data integrity.
Examples A malicious actor can add carefully designed perturbation to the input data to mislead the AI system
Severity High to Critical
Common Consequence Affects the integrity of the model’s output without manipulating the model itself.
Relationships None
Modes of Introduction Inference
Potential Mitigation Adversarial Detection & Input Reconstruction
References https://ieeexplore.ieee.org/document/9010939 [12], https://openreview.net/pdf?id=SyJ7ClWCb [13]
TABLE II: Example 1- A potential AI-CWE entry: Inadequate Input Filtering
AI-CWE-ID AI-CWE-101
AI-CWE-Name Lack of Appropriate Privacy Safeguard
Description Privacy safeguards are deficient or absent, exposing the data to unauthorized access and manipulation due to inadequate encryption, anonymization, or access control configurations.
Examples The lack of appropriate privacy safeguards allows malicious actors to perform black-box attacks against the system to steal private information about the system
Severity Critical
Common Consequence Information Disclosure
Relationships None
Weakness Phase Inference, Training
Potential Mitigation Differential Privacy & Prediction Restriction
References https://ieeexplore.ieee.org/abstract/document/9745062, https://ieeexplore.ieee.org/document/10214061 [14], https://proceedings.neurips.cc/paper_files/paper/2023/hash/e9df36b21ff4ee211a8b71ee8b7e9f57-Abstrac.t-Conference.html [15]
TABLE III: Example 1- A potential AI-CWE entry: Lack of Privacy Safeguard
II-B Impact
Describes the repercussions of a vulnerability within the AI system, detailing how the flaw affects the functionality, security, and operational reliability. It evaluates both direct and indirect effects, such as system performance degradation, potential exploitation risks, and broader implications for system users and associated services.
II-C Severity Scores
The Common Vulnerability Scoring System (CVSS) is a widely used standard for assessing the severity of security vulnerabilities in software. However, in the context of AI systems, the unique characteristics of these technologies expose several gaps in the CVSS and other similar framework. Given the distinctive nature of AI vulnerabilities, a tailored approach to severity assessment is necessary. Below, we highlight the gaps in current methodologies like CVSS, discuss how to measure the severity of AI vulnerabilities, and propose viable directions.
There are several challenges in severity measurement of AI vulnerabilities.
II-C1 CVSS Base Metrics and AI
  • Exploitability Metrics (Attack Vector, Attack Complexity, Privileges Required, User Interaction): These metrics provide a fundamental understanding of how a vulnerability can be exploited. However, AI systems may introduce additional complexities not captured by these metrics, such as the subtle ways data manipulation can influence AI behavior.
  • Impact Metrics (Confidentiality, Integrity, Availability): While these metrics address direct impacts of vulnerabilities, AI systems might suffer from indirect effects such as decision-making errors or biases introduced through data poisoning, which are not directly related to data loss or system downtime, but affect the system’s operational integrity.
II-C2 Environmental Metrics and AI
These metrics allow adaptation to particular organizational infrastructures, which is crucial for AI systems as their deployment contexts can significantly influence vulnerability impacts. However, the dynamics of AI behavior, learning over time, and interaction with varying data streams might require even more dynamically adjustable metrics.
II-C3 Supplemental Metrics and AI
Safety, Automatable, Recovery, and Value Density metrics are designed to provide additional context. For AI systems, the safety metric is particularly crucial as AI decisions might directly impact human safety. The automatable metric is especially important in the context of AI, considering the potential for automated attacks exploiting AI vulnerabilities across multiple systems.
II-C4 Notable Gaps
Despite the comprehensive nature of CVSS, there are several notable gaps when applied to AI: (1) Dynamic Nature of AI: AI systems evolve, learning from new data, which might alter their vulnerability landscape over time. CVSS scores, being largely static once assigned, may not reflect the changing nature of AI vulnerabilities. (2) Indirect Impacts: AI vulnerabilities might not directly compromise data confidentiality, integrity, or availability but could degrade the quality of AI decisions, leading to erroneous outcomes. These impacts are not well captured by current CVSS metrics. (3) Data and Model Specific Attacks: Attacks on AI may involve data manipulation or model evasion techniques that do not have direct analogs in traditional software vulnerabilities, requiring new types of impact and exploitability metrics.
II-C5 Considerations
Adapting the CVSS to better fit the characteristics of AI vulnerabilities involves a deeper examination of how AI-specific risks can be systematically and dynamically assessed.
  • Dynamic Scoring: To address the dynamic nature of AI systems, where the model’s behavior can change over time due to learning processes, a dynamic scoring system could be implemented. This would involve periodic reassessments of vulnerabilities based on how the AI model evolves. One approach could be to use automated monitoring tools that track changes in the model’s behavior and reevaluate vulnerability scores at defined intervals or when significant changes occur in the model’s input data or learning parameters. Implementing dynamic scoring raises questions about the frequency and triggers for score updates, the stability of vulnerability management programs, and the computational overhead involved in continuously monitoring AI systems.
  • AI-Specific Metrics: Developing AI-specific metrics involves identifying and defining impacts and exploitabilities unique to AI systems. While measuring the impact of AI vulnerabilities on Confidentiality, Integrity, and Availability (CIA) is important, it is also necessary to develop a new set of metrics meaningful for AI software. For instance, metrics could measure susceptibility to data poisoning, resistance to model inversion attacks, and sensitivity to adversarial examples. Another area could be the robustness of the model under distributional shift, where the data’s statistical properties change over time, which is not currently addressed in traditional CVSS metrics. The main challenge is the technical complexity involved in quantifying such metrics, especially in a standardized manner that fits diverse AI applications across different industries.
This area remains open and requires further research within the community to develop a robust and accurate severity scoring system.
II-D Affected Software and Products
Specifies the software products, AI models, or systems that contain or are susceptible to the vulnerability. This may include versions, configurations, or specific components that are affected. This will require additional research to map a reported vulnerability to a unique set of AI products.
II-E Exploitability
Describes how a vulnerability can be exploited and reproduced. In this field, the following details are considered:
  • Technical Complexity: Describes the technical skills and knowledge required to exploit the vulnerability in the AI system. This might include the sophistication of the techniques that must be employed, such as crafting inputs that deceive AI models (e.g., adversarial examples) or exploiting specific weaknesses in the AI’s design.
  • Required Privilege Level: Indicates whether an attacker needs elevated privileges to exploit the vulnerability or if it can be accessed with no or minimal privileges. For AI systems, this could also include access to various properties such as model parameters, data, or the inference model.
  • Specific Actions Required: Details the steps that an attacker must take to successfully exploit the vulnerability.
II-F Mitigation
This section details strategies to address vulnerabilities in AI systems. It assesses vulnerabilities based on the specific AI model, data, and domain, providing tailored strategies for patching flaws and enhancing system resilience. This may include software patches, model updates, or data processing revisions. Additionally, broader system upgrades may be recommended to strengthen defenses against future threats, ensuring enhanced security and system integrity. Each entry in the mitigation section offers a comprehensive analysis of a technique, detailing its effects, advantages, and disadvantages, as well as the vulnerabilities it addresses. For example, the Adversarial Example Detection technique[16] serves as a countermeasure against adversarial machine learning attacks. TableIV presents the details of this mitigation technique.
Description The technique employs distance-based anomaly detection to identify adversarial examples using a subset of trusted data points. It divides these data points into two groups and trains a distance-based outlier detector for each group. During training, it establishes a pair of thresholds that are used to detect outliers. This approach helps in distinguishing between legitimate and adversarial inputs by analyzing the distances from these trusted points.
Effect Removes adversarial examples
Type Proactive (to be performed before or during training)
Tactic Adversarial Detection
Orientation Data
Target Weakness Inadequate Input Filtering
Target Attack Poisoning Attack
Pros Computationally efficient for large datasets and data with a large number of features
Cons Not effective against label flipping poisoning attacks
Reference https://arxiv.org/pdf/1802.03041
TABLE IV: An example of a Mitigation Technique
Description: The technique employs distance-based anomaly detection to identify adversarial examples using a subset of trusted data points. It divides these data points into two groups and trains a distance-based outlier detector for each group. During training, it establishes a pair of thresholds that are used to detect outliers. This approach helps distinguish between legitimate and adversarial inputs by analyzing distances from trusted points.
TABLE V: AI Vulnerability Details
Field Details
AI-CVE ID 2024-1234 (example ID)
Description The malicious actor first collects data similar to the target model’s training data to train multiple shadow models, which mimic the target model’s behavior. These models are then used to classify data points as ”in” (used in training) or ”out” based on their outputs. An attack model is subsequently trained on these classifications to learn the distinction between data used and not used in training. This attack model is finally applied to the target model’s outputs to infer whether specific data points were part of its training set, thus potentially exposing sensitive information if the attack is successful.
AI Model Detail Meta Data:
- Name: Google Prediction API & Amazon ML
- Type: Neural Networks (CNN)
- Version: 2017
- Dependencies: Torch 7
- Availability: Restricted Access
- Creator of the Model: Facebook (Meta)
- Certification and License: Google Cloud and AWS Certification
- Release Date: 05/22/2017
Architecture:
- Foundation and/or Additional Model: CNN
- Software and Hardware Requirements: Python3
- Dependencies: Torch7
Data:
- Data Source: NIST (MNIST)
- Collection: Handwritten digits collected from high school students
- Quality Assessment: Normalization, Similarizing training and test set
- Governance: Following 2007 Free Software Foundation, Inc. guidelines
- Annotation: High school students & United States Census Bureau
Consideration:
Ethical: Adheres to AWS and Google Cloud Code of Conduct
Environmental: Influenced by cloud machines in use
WeaknessType Lack of appropriate privacy safeguard
Root Cause(s) 1. Revealing entire prediction vector score
2. Insufficient coarsening of the prediction vector
Impact Expose individual data entries used in a model’s training set, leading to significant privacy violations and the risk of leaking confidential or personally identifiable information.
AI-Severity Score 9.0 / 10.0 (Critical). This vulnerability exposes confidential information, has Medium Exploitability, and affects a broad range of software and applications.
Affected Software & Products Google Cloud and Amazon Web Services
CPE ID(s) 2024/google/cloud/ModelV01, 2024/AWS/ModelV01
Exploitability The malicious actor needs to have (partial) access to the training data distribution, and should be able to query the target model for multiple times as required. The model also reveals the output vector of the prediction score.
Mitigation For this vulnerability, three mitigation strategies are advised: - Restricting the Prediction Vector to Top-K Classes: This technique involves limiting the model’s output to only the top-K probable classes for each prediction. By reducing the amount of detailed information available in the output, it becomes more challenging for an attack model to determine if a data point was part of the training set. - Coarsening the Precision of the Prediction Vector: This method reduces the granularity of the output probabilities, such as rounding or binning the values, which obscures the precision of how well a data point fits the model trained with the training data. This lack of precision hinders the attack model’s ability to make accurate membership inferences. - Differential Privacy: Adding controlled noise to the data or the model’s parameters during training, differential privacy ensures that the model’s outputs are approximately the same, regardless of the inclusion or exclusion of any individual data point. This approach not only protects against membership inference attacks but also helps comply with privacy regulations and builds user trust, especially in handling sensitive data.
References Membership Inference Attack and coarsening precision of the prediction scores[17], Differential Privacy[15, 18], Restricting Prediction to Top-k[19].
Report Date 03/25/2024
Reported by R. Shokri et al.[18]
Vendor Google Cloud (Alphabet), AWS (Amazon)
Status Resolved with defensive techniques
III AI Vulnerability Disclosure Process
The AI Vulnerability Disclosure Process should be tailored to address the unique characteristics and potential impacts of vulnerabilities within artificial intelligence systems. The process can leverage the best practices of software vulnerability disclosure processes to ensure efficient, responsible, and secure handling of vulnerabilities from discovery to resolution. This requires establishing organizational policies, assigning specific roles within the organization for handling disclosures, including a dedicated security team responsible for initial assessments, communications, and coordination with external researchers. Additionally, this requires conducting a preliminary analysis to verify the reported vulnerability. This may involve replicating the issue or reviewing the AI system’s code and behavior under controlled conditions. Within the disclosure process, the potential impact of the vulnerability on the AI system and its users needs to be evaluated. Factors such as the type of data at risk, the severity of the impact on system functionality, and the likelihood of exploitation need to be considered. Additionally, the criteria for responsible disclosure of AI vulnerabilities need to be developed.
Existing vulnerability management best practices can be adopted to support collecting, analyzing and determining the minimum elements. For instance, there will be a need for forming AI CVE Naming Authorities (CNAs), these will be organizations, vendors, researchers, Bug Bounty providers that are authorized to assign CVE IDs to vulnerabilities within their scope, facilitating standardized security vulnerability reporting and management. These entities will intake vulnerability reports and assign AI-CVE IDs and determine each of the minimum elements such as #1.AI-CVE ID, #7. Affected Software and Products and #8. CPE ID, #12. References, #13. Report Date will be determined through CNAs.
Existing sources such as AI/ML ATT&CK Framework[11, 20] can be utilizes to determine information required for the Minimum Elements such as #3. Weakness Type, #4. Root Cause, #5. Impact, and #10. Mitigation. Figure1 represents an overview of AI/ML ATT&CK Framework.
Figure 1: An overview of AI/ML ATT&CK Framework[11].
IV AI Bill of Materials to Support Minimum Elements
The AI Bill of Materials (AIBoM) is essential for providing deep insights into the origins and makeup of AI systems, thereby enhancing the traceability of components and the reproducibility of results. It plays a crucial role in identifying potential vulnerabilities and ensuring that ethical standards are maintained throughout the lifecycle of AI products. While there are established resources and guidelines for creating SBOM[21, 22], there is a notable absence of a standardized framework for AIBOM, highlighting a critical gap in the management of AI components. To the best of our knowledge, currently, there is no specific and comprehensive definition of AIBOM in the literature. Therefore, in this work, we propose AIBOM, a framework designed to fill this gap by providing structured guidelines and tools for documenting and managing diverse elements of AI systems. The proposed AIBOM will be used in AIVD to systematically catalog and assess vulnerabilities across AI models, enabling a more robust defense mechanism against potential threats and enhancing the overall security posture of AI deployments.
AIBOM consists of several components [23, 24], as shown in Figure2 which are detailed as follows:
Figure 2: An overview of the proposed AIBOM
IV-1 Meta Data
This component is responsible for providing a comprehensive overview of the AI model’s identity and background, ensuring clear documentation of its core characteristics.
  • Name: The name of the AI model.
  • Type: Type of the core model (type of the used machine learning or deep learning algorithm)
  • Version information
  • Dependencies: Software or other models that are prerequisites.
  • Availability: Open Source, Restricted Access, Private
  • Creator of the model
  • Certification such as official approval or compliance with industry standards.
  • Release Date
  • License that is legal permissions for use, modification, and distribution of the model.
IV-2 Model
: Details of the architecture and the model consisting of:
  • Foundation Model: Details about the underlying model used as a foundation.
  • Additional Model: Information about any supplementary models integrated with the base model. to the Foundation Model
  • Software Requirements
  • Hardware Requirements
  • Dependencies: Third party libraries and packages required for this model.
IV-3 Data
: Details of the Data used in this model including:
  • Data Source: Description of where and how data is sourced.
  • Collection Method: How the data is gathered.
  • Quality Assessment: Processes to ensure and evaluate data quality.
  • Governance: Policies and procedures managing data access, usage, and compliance.
  • Annotation Process: Details about labeling and annotation methods.
IV-4 Consideration
  • Ethical: Adherence to ethical guidelines and societal impact assessments.
  • Environmental: Analysis of energy consumption and carbon footprint.
IV-5 Usage
  • Intended Usage that is the primary application
  • Out of Scope & Malicious Usage that are scenarios that are not supported or are potentially harmful
V Motivating Examples
In this section, we present a real-world example that demonstrates how an AI vulnerability is identified, modeled, and managed within our AIVD framework. This example serves to illustrate the practical application of the AIVD in addressing a specific, documented vulnerability, highlighting the framework’s effectiveness in enhancing the reliability of AI technologies. Through this case study, we aim to showcase the capabilities of the AIVD in responding to actual threats faced by AI systems in various industries.
V-1 A Real-World AI Vulnerability
In this case, as shown in FigureV, we detail a vulnerability which malicious actors can infer the presence of individual data points in a model’s training set, which allows the malicious actor to deploy a Membership Attack, potentially violating data privacy and exposing sensitive information.
VI Moving Forward
We need to discuss the gaps, for instance “severity” can be part of AIVD, but how do we measure the severity of AI incidents?
As we aim to develop trustworthy and reliable AI systems, it is crucial to establish robust mechanisms for managing AI vulnerabilities. This section examines the urgent challenges and significant gaps that must be addressed to enhance the reliability and security of AI deployments, paving the way for more resilient AI frameworks.
  • Challenge #1: Complex Vulnerability Profiles: Traditional vulnerability databases like CVE and CWE are not equipped to fully capture the complexities of AI-specific issues. AI systems exhibit unique vulnerabilities due to their reliance on data quality, algorithmic complexity, and adaptive behaviors that standard vulnerability descriptors may not adequately address.
  • Challenge #2: Dynamic and Non-Deterministic Nature: The dynamic and often non-deterministic nature of AI systems complicates vulnerability identification. Traditional systems have static and predictable outputs, whereas AI outputs can change based on new data or algorithmic adjustments, making consistent vulnerability assessment challenging.
  • Challenge #3: Severity and Impact Assessment: Assessing the severity of vulnerabilities in AI systems poses significant challenges due to the variable contexts in which these systems operate. For example, the same vulnerability may have minor consequences in one scenario while causing catastrophic failures in another, depending on how the AI system is integrated and its critical role in decision-making.
  • Challenge #4: Interdependencies: AI systems often depend heavily on external data sources and dynamic learning algorithms, making them susceptible to vulnerabilities introduced through these dependencies. Identifying and tracing these vulnerabilities require a deep understanding of the interdependencies and the potential cascade effects.
  • Challenge #5: Data and Privacy Concerns: Many AI systems handle sensitive or personal data, which raises significant privacy concerns if vulnerabilities are exploited. This aspect necessitates not only tracking vulnerabilities but also ensuring compliance with data protection regulations.
  • Challenge #6: Update and Patch Management: Given the rapid development cycles of AI technologies and their applications, ensuring timely updates and patch management is a significant challenge. The continuous learning aspect of many AI systems may necessitate ongoing vulnerability assessments and updates, complicating traditional patch management strategies.
  • Challenge #7: Ethical and Bias Implications: AI systems can perpetuate or exacerbate biases present in their training data. Addressing these vulnerabilities requires not only technical solutions but also ethical considerations, which are currently not well-represented in existing vulnerability databases.
VII Conclusions
In this paper, we introduce the AIVD framework, designed to report and evaluate vulnerabilities within AI systems. We define the Minimum Elements necessary to accurately describe an AI vulnerability within this framework. Additionally, we propose AIBOM to describe and document the components and dependencies of AI systems in detail. This paper also provides examples of capturing and characterizing AI weaknesses and vulnerabilities and highlights challenges in effectively managing AI vulnerabilities. It underscores the importance of a systematic approach for identifying and mitigating potential threats in AI deployments.
VIII Acknowledgements
We thank the reviewers for their valuable comments.